# HashAttention: Semantic Sparsity for Faster Inference

### v1

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Quality of Llama-3.1-8B (Average of PassageRet, TriviaQA, MultiFieldQA and HotpotQA) for different sparse attention methods at different sparsity budgets with 32 bits per token auxiliary memory. At 32 $\times$ sparsity, HashAttention has only a average point difference of 0.52 points.
@@
The ability to refer to long contexts efficiently is crucial for modern AI applications, from processing lengthy documents to engaging in extended conversations (Touvron et al., 2023; Achiam et al., 2023; Liu et al., 2024b). It is commonplace for LLM models to preprocess and store huge amounts of text in the form of KV Cache, which is later used to process various prompts. The Scaled Dot Product Attention (SDPA), fundamental to the transformer architecture that has driven the Generative AI revolution (Vaswani, 2017; Brown et al., 2020), does not scale well with context length. Processing small prompts or generating a single token requires SDPA to access the entire context KV Cache, which can be hundreds of GB in size. For instance, a KV Cache of 256K tokens is 128GB in size (0.5 MB per token). Enabling sparsity in attention, where only a subset of tokens is used in each attention step, can significantly reduce the computational and memory burden in LLMs.
@@
Sparsity naturally arises in SDPA. Due to the softmax kernel, only a few tokens significantly contribute to the final attention computation(Bricken & Pehlevan, 2021; Deng et al., 2024). Efficiently identifying these pivotal tokens provides a pathway to achieving efficient attention. Various approaches to identifying pivotal tokens have been explored in the literature. Heuristic-based methods, such as fixed sparsity patterns(Xiao et al., 2023), ignore the dynamic nature of contextual sparsity in attention, resulting in suboptimal attention quality. KV cache discard strategies, such as those proposed in (Liu et al., 2024c; Zhang et al., 2023; Li et al., 2024a), identify the global importance of tokens per attention head, incorporating some degree of dynamic sparsity. These methods discard tokens based on their observed importance in context prefixes. However, token importance is dynamic, and tokens deemed unimportant in prefixes can become critical for future inference, leading these methods to fail in certain scenarios (Xiao et al., 2024; Tang et al., 2024). Some approaches to truly dynamic sparsity have emerged(Xiao et al., 2024; Yang et al., 2024; Tang et al., 2024). However, since these methods rely on heuristics to reduce the computation of pivotal tokens, they often exhibit low recall.
In this paper, we take a principled approach to identifying pivotal tokens. Given a query and a large set of key-value pairs, we frame the task of retrieving important tokens as a recommendation problem(Zhang et al., 2021). This approach allows us to leverage the extensive literature on information retrieval (IR)(Schütze et al., 2008) to devise an efficient and GPU-friendly solution. One of the key approaches in information retrieval is to learn an embedding space where the semantic similarity between an item (e.g., an attention key-value pair) and a query (e.g., an attention query) is reflected in the distance between their embeddings. HashAttention embeds the key-value pairs and queries in Hamming space using learned functions. Given a query, the nearest key-value embeddings are used to identify pivotal tokens. The binary embeddings for the key-value cache can be packed into integers and stored efficiently using minimal memory. Furthermore, HashAttention identifies pivotal tokens using bit-wise operations, making the retrieval process fast.
We demonstrate that HashAttention outperforms all relevant baselines in terms of quality per unit token budget. HashAttention can significantly reduce the number of tokens required to match the quality of the full model. Figure 1 shows average quality over a random subset of datasets from LongBench(Bai et al., 2024b). HashAttention can almost match the full model quality with 32×32$\times$32 × compression. To achieve similar quality competing methods can only afford a 8×8$\times$8 × compression. All methods for sparse attention, including HashAttention, differ only in their algorithms for computing pivotal tokens, while sharing the same underlying routine for computing sparse attention with the selected tokens. This routine is the bottleneck for overall latency. As a result, achieving superior recall enables models to run with fewer tokens, significantly improving overall efficiency. At 32×32$\times$32 × compression, HashAttention is upto 6×6$\times$6 × faster than LightLLM(LightLLM, 2024) and upto 4.5×4.5$\times$4.5 × faster than Gpt-Fast(GPTFast, 2024).
Furthermore, the additional costs of HashAttention are much lower than those of its competing approaches. First, memory utilization during LLM inference is a significant bottleneck, particularly due to the large sizes of KV caches, making memory a critical resource. The memory footprint of HashAttention is as small as one unsigned integer per token per attention head (PTPA) , compared to, for example, 16 channels PTPA in Double Sparsity or Quest. Second, the per-token computation in HashAttention requires bit operations, which are faster than Multiply-Add operations used by other approaches.
To summarize, we make the following contributions,
  • We propose HashAttention to learn the semantic space of KV pairs and queries via learnable hash functions. During inference, pivotal tokens are identified based on their proximity to the query in this semantic space.
  • We provide a lightweight post-training recipe for adapting HashAttention to pretrained LLM models without modifying the LLM parameters.
  • HashAttention significantly improves the recall of pivotal tokens compared to baselines, reducing the number of tokens required in attention computations. Specifically, HashAttention requires 4×4$\times$4 × fewer tokens than sparse-attention baselines.
  • HashAttention can achieve 32×32$\times$32 × sparsity with minimal loss of quality and improvement in latency of upto 6×6$\times$6 × over LightLLM and 4.5×4.5$\times$4.5 × over GPT-Fast.
@@ 2.2 Post-training Sparse Attention
Various approaches exist to sparsify the attention computation of pretrained models. These can be classified as follows:
@@
Approaches such as StreamingLLM (Xiao et al., 2023) adopt a fixed sparsity pattern to approximate attention based on the observed importance of attention sinks (e.g., the first few tokens) and local window tokens. However, subsequent studies (Zhang et al., 2023; Xiao et al., 2024) have convincingly demonstrated the dynamic nature of sparsity and the resulting quality degradation when using fixed sparsity.
@@
To manage long contexts, methods such as H2O(Zhang et al., 2023), ScissorHands(Liu et al., 2024c), FastGen(Ge et al., 2023), and SnapKV(Li et al., 2024a) discard tokens based on specific heuristics. However, once discarded, these tokens are no longer available for future generations. This limitation is particularly problematic in scenarios such as multi-turn chat or multiple QA sessions on a fixed document, where it is essential to access different parts of the document in each turn or question.
@@
Attention scores are one of the critical components that determine the contribution of a token to the output of the attention mechanism. Identifying the top-k tokens with the highest attention scores requires $O(nd)$ computation, as it involves the dot product of the query with each token in the KV cache, where $n$ n is the number of tokens in the KV cache and $d$ is the dimensionality.
Double Sparsity(Yang et al., 2024) reduces this computational cost by selecting fewer channels to estimate dot products, which are then used to identify the top-k tokens. The channels are chosen based on offline calibration of channel norms. InfLLM(Xiao et al., 2024) and Quest(Tang et al., 2024) reduce computation along the $n$-dimension by creating page-level representations. These methods include or exclude all tokens on a page at once. While these approaches are effective, they often fail to provide high recall for the top-k tokens with respect to attention scores. Additionally, it is important to note that the absolute importance of a token depends not only on its attention score but also on the norm of its value vector, a factor completely ignored by these methods. (see lemma 4.1).
@@
Recently, RetrievalAttention(Liu et al., 2024a) proposed using a graph-based nearest neighbor algorithm to identify the top-k tokens with the maximum inner product. RetrievalAttention offloads the top-k computation to the CPU due to the sparse computations involved with graphs, leading to additional latency. SqueezeAttention(Hooper et al., 2024), a concurrent work to ours, proposed solving the top-k problem efficiently by clustering keys. Both the methods ignore the contribution of value vectors in determining top-k and it is non-trivial to include this information.
2.3 Efficient Attention from scratch
Another line of work aimed at improving the efficiency of attention focuses on designing mechanisms to avoid quadratic computation. Notable efforts include linear attention methods based on Random Fourier features, such as Performerc(Choromanski et al., 2020), RFA(Peng et al., 2021), and LARA(Zheng et al., 2022). These approaches approximate the softmax, in expectation, using random features to compute attention weights. Linformer(Wang et al., 2020) reduces attention complexity by using low-rank approximations, and sketching long-context token embeddings along the sequence dimension. This method is inspired by the Johnson–Lindenstrauss Lemma(Johnson et al., 1986) for dimensionality reduction. Reformer(Kitaev et al., 2020) combines locality-sensitive hashing (LSH) tables with attention to reduce complexity through space partitioning. While these approaches have garnered significant interest in the research community, empirical evaluations have shown that Scaled Dot Product Attention (SDPA) demonstrates the most favorable scaling performance – i.e., as model sizes and training compute increase, the capabilities increase commensurately.
@@ 3 Background
Figure 2: The general recipe of a sparse attention. A cheap approximation scores the key-value pairs for determining top-k. Top-k tokens are then used for attention computation.
@@ 3.1 Scaled Dot Product Attention (SDPA)
The Scaled Dot Product Attention has remained the most promising attention and has stood the test of time. The computation of SDPA for key and value embeddings, $\mathbf{K,V}:n_{1}\times d$, and $\mathbf{Q}:n_{2}\times d$ can be written as,
@@
3.2 Recommendation
The problem in recommendation can be abstracted as follows. Given a set of items $\mathcal{I}$ and a user $u\in\mathcal{U}$, the recommendation aims to select a small subset of $\mathcal{I}$ that is relevant to $u$. The historical relevance of items for users is captured in an interaction matrix, which is $|\mathcal{U}|\times|\mathcal{I}|$ matrix. Traditionally, the learning recommendation model has been cast as a matrix factorization of the interaction matrix. The two matrices, thus obtained, provide us with embeddings of items and users that can be used for inference. Subsequently, auxiliary information such as description, profile, interaction history, etc., and deep learning models are used to obtain richer user and item representations. In this embedding space, given a user embedding, items that have the highest inner product are the ones most relevant to items. Alternatively, one can learn the embeddings such that relevant items lie close to user embeddings in terms of $l_{p}$ distance. To improve the efficiency of relevant item retrieval, we often use approximate near-neighbor algorithms and data structures built on the top of the embedding space.
Sparsifying attention is a recommendation problem with key-value pairs as items and queries as a user. We want to select key-value pairs that minimize the error of output embedding w.r.t full attention for a particular query.
@@ 4 HashAttention
Figure 3: HashAttention working: The key-value pairs are mapped to bit signatures via learned mapping function $\phi_{kv}$. The query is mapped to bit signature via function $\phi_{q}$. The tokens closest to query signature are chosen as candidates for attention computation.
4.1 General Setting of Sparse Attention
The general recipe of sparse attention can be viewed as a combination of two subroutines (1) SPARSIFY (2) SPARSE-ATT
SPARSIFY ($\mathcal{K},\mathcal{V},\mathbf{q},\mathbf{k}$): Given a query $\mathbf{q}$ and a set of tokens $\mathcal{K},\mathcal{V}$, SPARSIFY returns $k$ tokens with highest estimated importance from $\mathcal{K}$.
SPARSE-ATT ($\mathcal{K},\mathcal{V},\mathbf{q},\mathcal{I}$): Given a set of keys $\mathcal{K}$, associated values $\mathcal{V}$, a query $\mathbf{q}$ and set of indices $\mathcal{I}$ identifying tokens to be used, SPARSE-ATT computes the attention only using tokens indicated by $\mathcal{I}$ as follows,
@@ 4.1 General Setting of Sparse Attention
where $\mathbf{v}_{i}$, $\mathbf{k}_{i}$ refer to $i^{th}$ elements of $\mathcal{V}$ and $\mathcal{K}$ respectively, and $\mathbf{1}$ is the indicator function.
The SPARSIFY function can be further decomposed into two steps SCORE and TOPK . SCORE ($\mathcal{K},\mathcal{V},\mathbf{q}$) assigns a query-aware score to each token in $\mathcal{K}$. TOPK picks the top tokens with the highest scores assigned by SCORE . In the text, we will overload the function SCORE ($\mathbf{k},\mathbf{v},\mathbf{q}$) to denote the score assigned to a single key for methods that work on each key individually.
HashAttention along with various other previous methods fit this structure. The ingenuity of different methods lies in the SCORE function. The ordering established by SCORE needs to represent the importance of keys accurately. The quality of SCORE can be measured in terms of its $\textrm{recall}(n,k)$ defined as follows. Let $\mathcal{I}_{n}$ be the set of top $n$ indices chosen by SCORE function and $\mathcal{I}_{true,k}$ be the set of $k$ tokens with the highest true attention scores. Then the recall is
$\textrm{recall}(\mathcal{I}_{n},\mathcal{I}_{true,k})=\frac{|\mathcal{I}_{n}\cap\mathcal{I}_{true,k}|}{|\mathcal{I}_{true,k}|}$ (2)
Following the SCORE routine, most methods have identical routines of TOPK and SPARSE-ATT . As illustrated in experiments, most time in sparse attention is spent in TOPK and SPARSE-ATT routines. Thus we should build a high recall SCORE function with reasonably small latency.
SCORE function has three important aspects.
@@
    Auxiliary memory of SCORE : Various methods such as Double Sparsity(Yang et al., 2024), Quest(Tang et al., 2024), and InfLLM(Xiao et al., 2023) including HashAttention use meta representations of keys/tokens that are cached for efficient score computation. Since memory is an important bottleneck with huge sizes of KVCache, the amount of memory used for meta-information is an important consideration.
@@
    Latency of SCORE is another important factor. Since SCORE is generally not a latency bottleneck, we have some scope to be lenient in using more computation to obtain higher-quality retrieved tokens
@@
    Quality of SCORE : A SCORE with higher recall, implies we can effectively use fewer tokens in SPARSE-ATT which will further improve the overall efficiency of the attention procedure.
4.2 SCORE function for HashAttention
Identifying the top tokens for a query is analogous to user-item recommendation problem – which has been extensively studied in Information Retrieval. HashAttention uses two learnable mapping functions, $\phi_{kv},\phi_{q}:R^{d}\rightarrow[0,1]^{b}$ to lift key-value and queries from $R^{d}$ to hamming space of dimension $b$. In this space, we can efficiently compare a given query to the keys using hamming distance $\mathcal{H}$.The SCORE function is defined by

### v2

@@ Abstract
Leveraging long contexts is crucial for advanced AI systems, but attention computation poses a scalability challenge. While scaled dot-product attention (SDPA) exhibits token sparsity, i.e. only a few pivotal tokens significantly contribute to output, exploiting this sparsity remains challenging. Existing methods either suffer from quality degradation or require substantial additional resources. We show that identifying pivotal tokens is a Maximum Inner Product Search (MIPS) problem. However, existing MIPS solutions are not well-suited for SDPA, as they are not GPU-friendly and often underperform due to the separated query and key distributions. This paper introduces HashAttention, framing pivotal token identification as a recommendation problem. Given a query, HashAttention encodes keys and queries in Hamming space, capturing the required semantic similarity, using learned mapping functions. HashAttention efficiently identifies pivotal tokens for a given query using bitwise operations and computes attention using only these tokens, improving the overall attention efficiency. Trained on generic data, HashAttention reduces tokens used by up to 16×16$\times$16 × with minimal quality loss, requiring only 32 bits of auxiliary memory per token. Sparsity can be further improved to 32×32$\times$32 × through task-specific fine-tuning. On A100 GPU, at 32×32$\times$32 × sparsity, incorporating HashAttention reduces attention latency by up to 4.3×4.3$\times$4.3 × in GPT-FAST and 2.54×2.54$\times$2.54 × in FlashDecode, and achieves up to 3.12×3.12$\times$3.12 × higher throughput for GPT-FAST.
@@ 1 Introduction
Figure 1: Quality vs. Throughput (on A100 GPU) vs. Auxiliary memory at different sparsity rates for Double Sparsity (16 channels, 2bit and 4 bit quantization) and HashAttention implemented in GPT-FAST (batch=1) framework on RULER@32K using Llama-3.1-8B-Instruct. The impact of auxiliary memory is excluded from throughput computation due to the relatively small scale of the problem (batch=1 and 32K context length). Also, throughput is overestimated for Double sparsity by excluding quantization/dequantization in throughput computation.
@@
The ability to refer to long contexts efficiently is crucial for modern AI applications, from processing lengthy documents to engaging in extended conversations (Touvron et al., 2023; Achiam et al., 2023; Liu et al., 2024b). It is commonplace for LLM models to preprocess and store huge amounts of text in the form of KV Cache, which is later used to process various prompts. However, the Scaled Dot Product Attention (SDPA), which is fundamental to the transformer architecture that has driven the Generative AI revolution (Vaswani, 2017; Brown et al., 2020), does not scale well with context length. Generating a single token requires SDPA to access the entire context KV Cache, which can be hundreds of GB. For instance, a KV Cache of 512K tokens is 64GB (0.125 MB per token) for the LLama3.1-8B model.
@@
Luckily, sparsity naturally arises in SDPA. Due to the softmax kernel, only a few pivotal tokens significantly contribute to the final attention computation(Bricken & Pehlevan, 2021; Deng et al., 2024). Efficiently identifying these pivotal tokens provides a pathway to achieving efficient attention. Various approaches to identifying pivotal tokens have been explored in the literature. Heuristic-based methods, such as fixed sparsity patterns(Xiao et al., 2023), ignore the dynamic nature of contextual sparsity in attention, resulting in suboptimal attention quality. KV cache discard strategies, such as those proposed in (Liu et al., 2024c; Zhang et al., 2023; Li et al., 2024a), identify the global importance of tokens per attention head, incorporating some degree of dynamic sparsity. These methods discard tokens based on their observed importance in context prefixes. However, token importance is dynamic, and tokens deemed unimportant in prefixes can become critical for future inference, leading these methods to fail in certain scenarios (Xiao et al., 2024; Tang et al., 2024). Some approaches to truly dynamic sparsity have emerged(Xiao et al., 2024; Yang et al., 2024; Tang et al., 2024). However, firstly these methods rely on heuristics to reduce the computation of pivotal tokens, leading to low recall rates. Secondly, improving recall with these methods needs substantial auxiliary memory.
In this paper, we take a principled approach to identifying pivotal tokens. Given a query and a large set of key-value pairs, we frame the task of retrieving important tokens as a recommendation problem (Zhang et al., 2021). We show that identifying pivotal tokens under the independence of value vectors is exactly a Maximum Inner Product Search(MIPS) problem. Existing solutions to MIPS are unsuitable for pivotal token identification due to two reasons (1) SOTA solutions to MIPS are based on graph data structures which are GPU unfriendly. (2) Query and Key distributions are separated in Attention causing off-the-shelf solutions to fail (Liu et al., 2024a). We observe that MIPS can be reduced to a cosine-similarity search (Neyshabur & Srebro, 2015; Shrivastava & Li, 2014) which can be further approximated using bit signatures (Gionis et al., 1999), or embeddings in Hamming space.
HashAttention uses learned independent mappings to embed the key-value pairs and queries in a Hamming space. These mappings are trained on the actual key-value and query data from the LLM model to encode inner product-based similarity into bit signatures. Given a query, the tokens with key-value signatures closest to the query signature in Hamming space are identified as pivotal tokens. These signatures can be packed in integers enabling their efficient storage and processing. HashAttention is GPU friendly and resolves the issue of separated distribution of queries and keys using learned independent mappings.
An instance of throughput improvement for GPT-FAST, on RULER@32K benchmark with Llama-3.18B-Instruct is shown in Figure 1. At 32K context length, HashAttention improves the throughput 1.3×1.3$\times$1.3 × (4×4$\times$4 × sparsity) at quality drop of $0.8$ points and 1.7×1.7$\times$1.7 × (8×8$\times$8 × sparsity) with quality drop of $1.7$ points using auxiliary memory of 32 bits per token per attention head (PTPA). At auxiliary memory of even 64 bits PTPA, Double Sparsity(DS), a SOTA sparse attention, fails to preserve model quality and improve throughput. Broader experiments reveal more potential for sparsity and throughput improvements.
HashAttention, trained with a generic dataset (i.e. independent of downstream tasks) can give up to 16×16$\times$16 × reduction in KV cache usage while preserving the model quality. For instance, with 16×16$\times$16 × sparsity, HashAttention maintains model quality within a drop of 0.78 points on LongBench(Bai et al., 2024b) and 1.13 points on RULER@16K using Llama-3.1-8B-Instruct model. Our ablations show that sparsity can be further improved if HashAttention is finetuned on task-specific data. For instance, on some Longbench datasets with Llama-3.1-8B-Instruct and Mistral-7B-v.03-Instruct, we can go up to 32×32$\times$32 × sparsity without loss of model quality.
Incorporating HashAttention in inference systems can improve the latency and throughput of inference. For instance, HashAttention-32$\times$ improves latency of GPT-FAST(GPTFast, 2024) attention upto 4.3×4.3$\times$4.3 × and that of FlashDecode(Dao, 2023) upto 2.54×2.54$\times$2.54 ×. Additionally, end-to-end model throughput for GPT-FAST, tested with Llama3.1-8B model, can be improved up to 3.12×3.12$\times$3.12 × with HashAttention-32$\times$. Furthermore, HashAttention only uses 32 bits PTPA auxiliary memory. To achieve the same quality, other sparse-attentions either need around 4×4$\times$4 × more pivotal tokens (i.e. 4×4$\times$4 × less sparsity) at 32 PTPA or need more auxiliary memory to operate at the same sparsity. The code for HashAttention is here.  [^1] https://github.com/xAlg-ai/HashAttention-1.0
@@ 2.2 Post-training Sparse Attention
We classify the previously proposed approaches to sparsify the attention computation of pre-trained models as follows:
@@
Approaches such as StreamingLLM (Xiao et al., 2023) adopt a fixed sparsity pattern to approximate attention based on the observed importance of attention sinks (e.g., the first few tokens) and recent tokens. However, subsequent studies (Zhang et al., 2023; Xiao et al., 2024) have demonstrated the dynamic nature of sparsity and the resulting quality degradation when using fixed sparsity.
@@
Methods such as H2O(Zhang et al., 2023), ScissorHands(Liu et al., 2024c), FastGen(Ge et al., 2023), and SnapKV(Li et al., 2024a) discard tokens based on specific heuristics. However, once discarded, these tokens are no longer available for future generations. This limitation is particularly problematic in scenarios such as multi-turn chat or multiple QA sessions on a fixed document, where it is essential to access different parts of the document in each turn or question.
@@
Attention scores are one of the critical components that determine the contribution of a token to the output of the attention mechanism. Identifying the top-k tokens with the highest attention scores requires $O(nd)$ computation, as it involves the dot product of the query with each token in the KV cache, where $n$ is the number of tokens in the KV cache and $d$ is the dimensionality.
Double Sparsity (Yang et al., 2024) reduces this computational cost by selecting fewer channels to estimate dot products, which are then used to identify the top-k tokens. The channels are chosen based on offline calibration of channel norms. Loki (Singhania et al., 2024) reduces the dimension by using traditional dimensionality reduction techniques to obtain low-dimensional representation for keys. InfLLM (Xiao et al., 2024) and Quest (Tang et al., 2024) reduce computation along the $n$-dimension by creating page-level representations. These methods include or exclude all tokens on a page at once. While these approaches are effective, they often fail to provide high recall for the top-k tokens with respect to attention scores.
@@
Recently, RetrievalAttention (Liu et al., 2024a) proposed using a graph-based nearest neighbor algorithm to identify the top-k tokens with the maximum inner product. RetrievalAttention offloads the top-k computation to the CPU due to the sparse computations involved with graphs, leading to additional latency. SqueezedAttention (Hooper et al., 2024), a concurrent work with ours, proposed to solve the top-k problem efficiently by clustering keys. In contrast, we use succinct bit signatures from key-value and queries to identify pivotal tokens.
Estimating attention output using Hash tables
Learning-to-hash attention (Sun et al., 2021) proposes to learn hash table mappings to obtain efficient attention with sparse retrieval. They design a learning algorithm to obtain a balanced distribution of keys in the hash table buckets. MagicPig(Chen et al., 2024) noted that top-k attention is biased and can be problematic in cases when attention scores are relatively balanced. They propose importance sampling-based unbiased attention estimation. They use LSH tables to approximate the importance-sampling distribution. These methods can be potentially combined with HashAttention for further efficiency.
Figure 2: HashAttention working: The key-value pairs are mapped to bit signatures via learned mapping function $\phi_{kv}$. The query is mapped to bit signature via function $\phi_{q}$. The tokens closest to the query signature are chosen as candidates for attention computation.
@@ 3.1 Scaled Dot Product Attention (SDPA)
The computation of SDPA for key and value embeddings, $\mathbf{K,V}:n_{1}\times d$, and $\mathbf{Q}:n_{2}\times d$ can be written as,
@@
3.2 Learning Recommendation Model
The recommendation problem can be abstracted as: Given a set of items $\mathcal{I}$ and a user $u\in\mathcal{U}$, the recommendation aims to select a small subset of $\mathcal{I}$ that is relevant to $u$. The historical relevance of items for users is captured in an interaction matrix with shape $|\mathcal{U}|\times|\mathcal{I}|$. Traditionally, the learning of recommendation model has been cast as a matrix factorization of the interaction matrix. The two matrices, thus obtained, provide us with embeddings of items and users that can be used for inference. Subsequently, auxiliary information such as description, profile, interaction history, etc., and deep learning models were used to obtain richer user and item representations. Alternatively, one can learn the embeddings so that the embeddings of the relevant items lie close to the user embeddings in terms of $l_{p}$ distance. To improve the efficiency of relevant item retrieval, approximate near-neighbor algorithms and data structures are utilized on top of the embedding space. Sparsifying attention is a recommendation problem with key-value pairs as items and queries as a user. We want to select key-value pairs that minimize the error of output embedding w.r.t full attention for a particular query.
@@ 4 HashAttention
The general recipe of sparse attention can be viewed as a combination of three subroutines (1) SCORE (2) TOPK and (3) GATHER-ATT .
SCORE ($\mathcal{K},\mathcal{V},\mathbf{q}$): Given a query $\mathbf{q}$ and a set of tokens $\mathcal{K},\mathcal{V}$, SCORE assigns a score to each token in ($\mathcal{K},\mathcal{V}$). TOPK is the standard top-k on these scores. GATHER-ATT ($\mathcal{K},\mathcal{V},\mathbf{q},\mathcal{I}$): Given a set of tokens $\mathcal{K},\mathcal{V}$, a query $\mathbf{q}$ and set of indices $\mathcal{I}$ identifying pivotal tokens, GATHER-ATT computes the attention only using tokens indicated by $\mathcal{I}$ as follows,
@@
$\;\;=\sum_{i=1}^{|\mathcal{K}|}\frac{\mathbf{1}(i\in\mathcal{I})\exp{(\langle\mathbf{q},\mathbf{k}_{i}\rangle)}\mathbf{v}_{i}}{\sum_{j=1}^{|\mathcal{K}|}\mathbf{1}(i\in\mathcal{I})\exp{(\langle\mathbf{q},\mathbf{k}_{j}\rangle)}}$
where $\mathbf{v}_{i}$, $\mathbf{k}_{i}$ refer to $i^{th}$ embeddings of $\mathcal{V}$ and $\mathcal{K}$ respectively, and $\mathbf{1}$ is the indicator function. In the text, we will overload the function SCORE ($\mathbf{k},\mathbf{v},\mathbf{q}$) to denote the score assigned to a single key for methods that work on each key individually.
Most sparse-attention methods can be viewed as a combination of the above three routines with differences in the SCORE function. Following the SCORE routine, sparse-attention methods can be implemented using identical routines for TOPK and GATHER-ATT . Furthermore, we find in experiments, that most time is spent in GATHER-ATT and TOPK routines. Thus we should devise a good quality SCORE function with reasonably small latency. The quality of SCORE can be measured in terms of its $\textrm{recall}(n,k)$ defined as follows. Let $\mathcal{I}_{n}$ be the set of top $n$ indices chosen by SCORE function and $\mathcal{I}_{true,k}$ be the set of $k$ true pivotal tokens. Then, the recall is computed as
$\textrm{recall}(\mathcal{I}_{n},\mathcal{I}_{true,k})=\frac{|\mathcal{I}_{n}\cap\mathcal{I}_{true,k}|}{|\mathcal{I}_{true,k}|}$
The SCORE function has three important aspects.
@@
    Auxiliary memory: Various methods such as Double Sparsity, Quest, and InfLLM including HashAttention use meta representations of keys/tokens that are cached for efficient score computation. Since memory is an important bottleneck with huge sizes of KVCache, the amount of memory used for meta-information is an important consideration.
@@
    Latency is another important factor. Since SCORE is generally not a latency bottleneck, one can devote more computation to obtain higher-quality retrieved tokens
@@
    Quality: A SCORE with higher recall, implies we can effectively use fewer tokens in GATHER-ATT which will further improve the overall efficiency of the attention procedure.
4.1 The SCORE function for HashAttention
Identifying the top tokens for a query is analogous to a user-item recommendation problem. HashAttention uses two learnable mapping functions, $\phi_{kv}:R^{2d}\rightarrow[0,1]^{b},\phi_{q}:R^{d}\rightarrow[0,1]^{b}$ to lift key-value and queries to hamming space of dimension $b$. In this space, we can efficiently compare a given query to the keys using hamming distance $\mathcal{H}$. The SCORE function is defined by